OpenStudy (anonymous):

6. Write the equation in slope-intercept form. What are the slope and y-intercept? -5x-12y=11 A. y=-5/12x+11/12; slope: -5/12; y-intercept: 11/12 B. y=5/12x-11/12; slope: 5/12; y-intercept: -11/12 C. y=-5/12x-11/12; slope: -5/12; y-intercept: -11/12 D. y=5/12x+11/12; slope: 5/12; y-intercept: 11/12

OpenStudy (anonymous):

-5x - 12y = 11 -12y = 5x + 11 y = -5/12x - 11/12 slope is -5/12 y int is -11/12

OpenStudy (anonymous):

any questions ?

OpenStudy (anonymous):

no i just always check my answers before summiting them thanks :)

OpenStudy (anonymous):

anytime :)